1) Find the value of (x+y)^2 if x^2 - y^2 = -12 and y-x = 2
2) Given that (x+y)^2 = 289 and x^2 + y^2 =157 , determine the value of (x-y)^2

Answer by mananth(12270)

1) Find the value of (x+y)^2
if x^2 - y^2 = -12
(x+y)(x-y)=-12..........1
and y-x = 2
x-y=-2..............2
substitute x-y in (1)
(x+y)(-2)=-12
(x+y)=-12/-2
(x+y)=6
(x+y)^2= 36
.............
2) Given that (x+y)^2 = 289
x^2+2xy+y^2=289
x^2+y^2=157
157+2xy=289
2xy=289-157
2xy=132
and x^2 + y^2 =157 ,
determine the value of (x-y)^2
x^2-2xy+y^2
(x^2+y^2)-2xy
x^2 + y^2 =157, 2xy = 132
157-132= 25
m.ananth@hotmail.ca